If Mark spends a day, he can produce 1 pound of tea or 1 tray of biscuits. Zach can produce 3 pounds of tea or 2 trays of biscuits, and Broden can produce 1 pound of tea or 3 trays of biscuits in a day.

What is Mark’s, Zach’s and Broden’s opportunity cost for producing a tray of biscuits?
Answer #1

opportunity cost is the amount forgone to produce one unit of good

Opportunity cost for Mark for 1 tray of biscuits = 1 pound of tea

Opportunity cost for Zach for 1 tray of biscuits = 3/2 = 1.5 pound of tea

Opportunity cost for Broden for 1 tray of biscuit = 1/3= 0.33 pounds of tea.